OpenStudy (winsteria):

Which answer describes the solution to each equation? 3x+5=1/5(15x+10) A. one solution B. no solution C. an infinite number of solutions

OpenStudy (winsteria):

@AlexandervonHumboldt2 @KamiBug @acxbox22

OpenStudy (acxbox22):

multiply by 5 on both sides so it is 5(3x+5)=15x+10 15x+25=15x+10 25=10 but we know 25 doesnt equal 10 so what do you think is the answer?

OpenStudy (winsteria):

B?

OpenStudy (winsteria):

Thank you though